How do I find which two triangles are similar? a) 48 degree and 69 degree b) 63 degree and 48 degree c) 63 degree and 78 degree
Answer by adamchapman(301)

All the angled inside a triangle add up to 180 degrees.
Let the angle we dont know in triangle (a) be called "A"
48+69+A=180
A=180-69-48=63 degrees
Now list all the anlges in triangle (a): 48, 63 and 69 degrees
Now do the same for triangle (b):
63+48+B=180
B=69 degrees
Now list all the anlges in triangle (b): 48, 63 and 69 degrees
Therefore triangles (a) and (b) are similar.
The unknown angle in triangle (c) is:
C=180-63-78=39 degrees
There are no 39 degree angles in triangles (a) and (b), so (c) is not similar to the other triangles.
I hope this helps
P.S. I am trying to start up my own homework help website. I would be extremely grateful if you would e-mail me some feedback on the help you received to adam.chapman@student.manchester.ac.uk